OpenStudy (anonymous):

As a homework I am supposed to integrate sinxcosx using the integration method of substitution. I thought taking sinx as the inner-function might be a good idea because the derivative of the inner-function the cosx is already here but what the outer function? Integration by parts seems more appropriate here...

razor99 (razor99):

srry i didt get it

OpenStudy (anonymous):

but then I do the integral of tdt after replacing sinx with t and cosx with dt and get that the integral of sinxcosx=0.5*x^2

OpenStudy (mimi_x3):

Woops, my bad. \[\large \int\limits sinxcosx \] Let u = \(cosx\) ; \(\frac{du}{-sinx} \)

razor99 (razor99):

could u tell the question in numbers

OpenStudy (mimi_x3):

You can use both or either sinx or cosx for the substitution btw.

OpenStudy (dumbcow):

integration by parts could work....substitution would be more popular and simpler because the 1 function is derivatives of other u = sin x du = cos x dx --> dx = du/cos x cos x cancels --> integral u du = u^2/2 = sin^2(x)/2

OpenStudy (anonymous):

thx now I got it